A company estimates that 0.5% of their products will fail after the original warranty period but within 2 years of the purchase, with a replacement cost of $150.

If they offer a 2 year extended warranty for $19, what is the company's expected value of each warranty sold?
1
Expert's answer
2020-06-30T18:37:11-0400

Solution:

let x be the outcomes are $19 and -($150-19)=-$131


p(x) be the probabilities of outcomes are 1-0.005=0.995 and 0.005 respectively


Expected value =\Sigma x*p(x)

Σx∗p(x)

=(19*0.995)+(-131*0.005)

=18.905-0.655

=18.25


Answer:$18.25